OpenStudy (anonymous):

Some students planned for a get-together. The budget for food was $500. Five of the students failed to come because of the distance, and therefore the cost of food for each member increased by $5. How many students attended the get-together?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

show work please?

OpenStudy (anonymous):

\[\frac{500}{x-5} - \frac{500}{x} = $ 5\]

OpenStudy (anonymous):

thank you